OpenStudy (anonymous):

MEdal given The total cost for 8 bracelets, including shipping was $54. The shipping charge was $6. Write an equation that models the cost of each bracelet.

OpenStudy (anonymous):

8b + 6 = 54 8b = 54 - 6 8b = 48 b = 6 check... 8 bracelets at $6 = 48 + 6 (shipping) = 54 -- correct

OpenStudy (anonymous):

8x+6=54

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

yep :) 8x+6=54 then solve for x

OpenStudy (anonymous):

x=6

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

okay thank you, i have one more

jigglypuff314 (jigglypuff314):

sure :)

OpenStudy (anonymous):

The total cost for 8 bracelets, including shipping was $54. The shipping charge was $6. Determine the cost for each bracelet. Show your work

jigglypuff314 (jigglypuff314):

? wouldn't that be the same question? equation: 8x + 6 = 54 cost: x = 6

OpenStudy (anonymous):

8x+6=54 8x=54-6 8x=48 x=48/8 x=6 the bracelet cost $6 is that right

jigglypuff314 (jigglypuff314):

correct! :)

OpenStudy (anonymous):

aha thank you, u think u could help me with a couple more?

jigglypuff314 (jigglypuff314):

sure! :)

OpenStudy (anonymous):

ok :)

jigglypuff314 (jigglypuff314):

solve it like you would for any other equation :) so add 8 to both sides then divide both sides by 6

OpenStudy (anonymous):

14y-8 <18?

OpenStudy (anonymous):

then divide?

jigglypuff314 (jigglypuff314):

hmm??? 6y - 8 < 10 +8 +8 6y < 18 then divide :)

OpenStudy (anonymous):

ohhhh got confused theree lol ok got it

OpenStudy (anonymous):

so 1y<3

jigglypuff314 (jigglypuff314):

yep :) so you get y <_ 3

OpenStudy (anonymous):

ahh ok thank u

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

The figures above are similar. Find the missing length. Show your work

jigglypuff314 (jigglypuff314):

solve for x when \[\frac{ 3 }{ 5 } = \frac{ x }{ 3 }\]

OpenStudy (anonymous):

1.8?

OpenStudy (anonymous):

well 9/5

jigglypuff314 (jigglypuff314):

correct :)

OpenStudy (anonymous):

Sharon earns $25 per item she sells plus a base salary of $100 per week. Write and solve an inequality to find how many items she must sell to earn at least $700 per week. Provide your conclusion as a complete sentence.

OpenStudy (anonymous):

:)

jigglypuff314 (jigglypuff314):

idk complete sentence >,< equation would be 25x + 100 >_ 700

OpenStudy (anonymous):

ok haha ill do that one last, how about

OpenStudy (anonymous):

18. Solve the inequality. Show your work. –6b > 42 or 4b > –4

jigglypuff314 (jigglypuff314):

–6b > 42 divide both sides by -6 and don't forget to flip the sign :) 4b > –4 divide both sides by 4

OpenStudy (anonymous):

okay :) // Solve the equation. Show your work. 8x – 12 = 4x + 24

jigglypuff314 (jigglypuff314):

subtract 4x from both sides add 12 to both sides divide both sides by 4

OpenStudy (anonymous):

okay wait i got stuck on #18

OpenStudy (anonymous):

:/

jigglypuff314 (jigglypuff314):

? what's #18?

OpenStudy (anonymous):

Solve the inequality. Show your work. –6b > 42 or 4b > –4 this one

jigglypuff314 (jigglypuff314):

–6b > 42 divide both sides by -6 and don't forget to flip the sign :) /-6 /-6 b < -7 4b > –4 divide both sides by 4 /4 /4 b > -1 so you get b < -7 or b > -1

OpenStudy (anonymous):

ohhh okay, thank u :)

OpenStudy (anonymous):

now im on this one// Solve the equation. Show your work. 8x – 12 = 4x + 24

jigglypuff314 (jigglypuff314):

subtract 4x from both sides add 12 to both sides divide both sides by 4

OpenStudy (anonymous):